OpenStudy (anonymous):

can someone show me how to do this?....given the equation x^2+5x+4, find the x-intercepts

OpenStudy (anonymous):

that's not an equation. i assume you mean y = x^2 + 5x + 4 you have to solve x^2 + 5x + 4 = 0 so (x + 4)(x + 1) = 0 x = -4, -1